OpenStudy (anonymous):

2.) For each sum find: a.) the number of terms, b.) the first term, c.) the last term, and d.) evaluate the sum. (4 pts each) A. 5 B. 8 E (3n+1) E (2n/3) n=1 n=1

OpenStudy (anonymous):

\[\sum_{n=1}^{5}\left( 3n+1 \right)\] no of terms=5 first term =3*1+1=4 last term=3*5+1=16 \[n=5,\sum=\frac{ n }{ 2 }\left( a+l \right)=\frac{ 5 }{ 2 }\left( 4+16 \right)=?\]

OpenStudy (anonymous):

50 ?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

So then for the next one? I can do the same. Can you help me set that one up?

OpenStudy (anonymous):

yes you should proceed in the same way,i will help.

OpenStudy (anonymous):

Ok, I'm going to attempt it, one second!

OpenStudy (anonymous):

I'm a little confused.. No. of terms= 8

OpenStudy (anonymous):

First term is? 2*3+1?

OpenStudy (anonymous):

\[\sum_{n=1}^{8}\frac{ 2n }{ 3 }\] no .of terms n=8 first term \[a=\frac{ 2*1 }{ 3 }=\frac{ 2 }{ 3 }\] last term \[l=\frac{ 2*8 }{ 3 }=\frac{ 16 }{ 3 }\] use the above given formula and find sum.

OpenStudy (anonymous):

Could you go into slightly more detail. n=8 E= 8/2(2/3+16/3)

OpenStudy (anonymous):

So the answer would be around 24?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

correct.